# Value-Penalized Auxiliary Control from Examples for Learning without Rewards or Demonstrations

### v1

@@ Abstract
Learning from examples of success is an appealing approach to reinforcement learning that eliminates many of the disadvantages of using hand-crafted reward functions or full expert-demonstration trajectories, both of which can be difficult to acquire, biased, or suboptimal. However, learning from examples alone dramatically increases the exploration challenge, especially for complex tasks. This work introduces value-penalized auxiliary control from examples (VPACE); we significantly improve exploration in example-based control by adding scheduled auxiliary control and examples of auxiliary tasks. Furthermore, we identify a value-calibration problem, where policy value estimates can exceed their theoretical limits based on successful data. We resolve this problem, which is exacerbated by learning auxiliary tasks, through the addition of an above-success-level value penalty. Across three simulated and one real robotic manipulation environment, and 21 different main tasks, we show that our approach substantially improves learning efficiency. Videos, code, and datasets are available at https://papers.starslab.ca/vpace.
Keywords: Inverse Reinforcement Learning, Multitask Learning, Manipulation
1 Introduction
Robotics presents a unique challenge to learning algorithms: feedback is often formulated as a manually-defined dense reward function, or demonstration trajectories of an expert completing the task, both of which can be difficult to acquire, suboptimal, or biased. Ensuring that expert trajectories are nearly optimal, or, alternatively, learning an optimal policy from mixed or suboptimal data, are both challenging, open problems in imitation learning [1]. Sparse reward functions are less biased [2], but require significant exploration and can be non-trivial to acquire.
@@ 1 Introduction
We consider another form of feedback—example states of completed tasks. Obtaining example states can be far less laborious than designing a reward function or gathering trajectories: practitioners can gather states from a distribution that represents a completed task without consideration of how the states are reached by the expert. Learning from example states, equivalently referred to as example-based control [3], can be inefficient, however. Similar to sparse rewards, the example states provide no information about which actions led to the goal state(s). In robotics applications, this inefficiency is particularly undesirable, since executing suboptimal policies is costly both in terms of potentially destructive environmental effects and time. In this work, we aim to address the following question:
Is it possible to learn policies efficiently given only example states of completed tasks?
To answer this question, we propose a new example-based control method, value-penalized auxiliary control from examples (VPACE).
Figure 1: VPACE in a standard reinforcement learning loop and multitask actor-critic update. Value penalization ensures that all example state buffers have higher value than all other states, while auxiliary control from examples allows us to learn from auxiliary task examples instead of main tasks alone.
Our approach is inspired by LfGP [4], a method that introduces the use of a scheduler and expert trajectories of auxiliary tasks to improve exploration. We build upon this idea to leverage only example states rather than expert trajectories. Our contributions are fourfold: (i) We find that the naïve application of scheduled auxiliary tasks to example-based control can result in poorly-calibrated value estimates, and we alleviate this problem by introducing value penalization based on the current expected value of the provided examples.
Figure 2: Average normalized results for all tasks studied in this work.
The combination and complimentary nature of these two additions results in a dramatic improvement in sample efficiency and performance across four environments with 19 simulated and two real tasks (see Fig. 2) compared with prior state-of-the-art example-based control (RCE, [3]) and imitation learning (DAC, [5], SQIL [6]) approaches. (ii) We situate our approach among these baselines, showing the effects of mixing value penalization and auxiliary control under various reward-modelling assumptions. (iii) We demonstrate that VPACE can be as sample-efficient as learning algorithms with other forms of feedback including expert trajectories and reward functions. (iv) Finally, we theoretically show that our value-penalized objective does not prevent an agent from learning the optimal policy.
2 Related Work
@@
Sparse rewards are a desirable form of feedback for learning unbiased, optimal policies in reinforcement learning (RL), but they can be difficult to obtain, and present an immense exploration challenge on long-horizon tasks [7]. Reward shaping [8] and dense rewards can help alleviate the exploration problem in robotics [9, 10, 11], but designing dense rewards is difficult for practitioners [12], and can lead to surprising, biased, and suboptimal policies. An alternative to manually-defined rewards is to perform inverse RL (IRL), in which a reward function is recovered from demonstrations, and a policy is learned either subsequently [13, 14, 15] or simultaneously, in a process known as adversarial imitation learning (AIL) [5, 6, 16, 17, 18]. AIL actor-critic approaches can be further divided into methods that learn both a value function and a separate reward model [16, 17] or methods that learn a value function only [5, 6, 18].
@@
Like dense rewards, full trajectory demonstrations can be hard to acquire, suboptimal, or biased. Unlike IRL/AIL, in example-based control (EBC), a learning agent is only provided distributions of single successful example states. Previous EBC approaches include using generative AIL (GAIL, [16]) directly (VICE, [19]), soft actor critic (SAC, [20]) with an additional mechanism for generating extra success examples (VICE-RAQ, [21]), learning the goal distribution as a reward function (DisCo RL, [22]), performing offline RL with conservative Q learning (CQL, [23]) and a learned reward function [24], and using SAC with a classifier-based reward (RCE, [3]).
@@
All EBC methods can naturally suffer from poor exploration, given that success examples are akin to sparse rewards. Hierarchical reinforcement learning (HRL) aims to leverage multiple levels of abstraction in long-horizon tasks [25], improving exploration in RL both theoretically [26, 27, 28] and empirically [29, 30]. Scheduled auxiliary control (SAC-X, [31]) combines a scheduler with semantically meaningful and simple auxiliary sparse rewards. SAC-X has also been extended to the domain of imitation learning with full trajectories (LfGP, [4, 32, 33]). Our approach builds on aspects of VICE-RAQ [21], SQIL [6], RCE [3], and LfGP [4], in that we use off-policy learning without a separate reward model to maximize efficiency.
@@
3 Example-Based Control with Value-Penalization and Auxiliary Tasks
@@
Our goal is to generate an agent, with as few environment interactions as possible, that can complete a task given final state examples of both the successfully completed task and a small set of reusable auxiliary tasks, with no known reward function or full-trajectory demonstrations. We begin by formally describing the problem setting for example-based control in Section 3.1. In Section 3.2, we describe how scheduled auxiliary tasks can be applied to example-based control. Finally, motivated by the increased exploration diversity of the multitask framework, we propose a new Q-estimation objective in Section 3.3 that leverages value penalization for improved learning stability.
3.1 Problem Setting
@@
A Markov decision process (MDP) is defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,\mathcal{P},\rho_{0},\gamma\rangle$, where the sets $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action space, $\mathcal{P}$ is the state-transition environment dynamics distribution, $\rho_{0}$ is the initial state distribution, $\gamma$ is the discount factor, and the true reward $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is unknown. Actions are sampled from a stochastic policy $\pi(a|s)$. The policy $\pi$ interacts with the environment to yield experience $\left(s_{t},a_{t},s_{t+1}\right)$, generated by $s_{0}\sim\rho_{0}(\cdot),a\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$. The gathered experience $\left(s_{t},a_{t},s_{t+1}\right)$ is then stored in a buffer $\mathcal{B}$, which may be used throughout learning. When referring to finite-horizon tasks, $t=T$ indicates the final timestep of a trajectory. For any variables $x_{t},x_{t+1}$, we may drop the subscripts and use $x,x^{\prime}$ instead when the context is clear.
@@
In this work, we focus on example-based control, a more difficult form of imitation learning where we are only given a finite set of example states $s^{*}\in\mathcal{B}^{*}$, where $\mathcal{B}^{*}\subseteq\mathcal{S}$ and $\lvert\mathcal{B}^{*}\rvert<\infty$, representing a completed task. The goal is to (i) leverage $\mathcal{B}^{*}$ and $\mathcal{B}$ to learn or define a state-conditional reward function $\hat{R}:\mathcal{S}\to\mathbb{R}$ that satisfies $\hat{R}(s^{*})\geq\hat{R}(s)$ for all $(s^{*},s)\in\mathcal{B}^{*}\times\mathcal{B}$, and (ii) learn a policy $\hat{\pi}$ that maximizes the expected return $\hat{\pi}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\hat{R}(s_{t})\right]$. For any policy $\pi$, we can define the value function and Q-function respectively to be: $V^{\pi}(s)=\mathbb{E}_{a\sim\pi}\left[Q^{\pi}(s,a)\right],$ (1) $Q^{\pi}(s,a)=\hat{R}(s)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}}\left[V^{\pi}(s^{\prime})\right],$ (2) corresponding to the return-to-go from state $s$ (and action $a$). Both the value function and the Q-function for any policy $\pi$ satisfy the above Bellman equations that can be used for reinforcement learning (RL), specifically by temporal difference (TD) algorithms.
One way to learn the reward function is through adversarial imitation learning (AIL)—the learned reward function $\hat{R}$ is derived from the minimax objective [5, 16, 17]:
$\mathcal{L}(D)=\mathbb{E}_{s\sim\mathcal{B}}\left[\log\left(1-D(s)\right)\right]+\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[\log\left(D(s^{*})\right)\right],$ (3)
where $D$ attempts to differentiate the occupancy measure between the state distributions induced by $\mathcal{B}^{*}$ and $\mathcal{B}$. The output of $D(s)$ is used to define $\hat{R}(s)$, which is then used for updating the Q-function with Eq. 2. In the single-task regime, this is algorithmically identical to state-only AIL approaches with full demonstrations [34, 35].
3.2 Learning a Multitask Agent from Examples
Example-based control presents a challenging exploration problem for complex tasks. We alleviate this problem by adapting learning from guided play (LfGP) [4], an approach for improving exploration by learning from auxiliary-task expert data, in addition to main task data. Auxiliary tasks are selected to have semantic meaning (e.g. reach, lift). Individual task definitions, and sometimes the auxiliary example data themselves, are reusable between main tasks.
LfGP augments an MDP $\mathcal{M}$ to contain auxiliary tasks, where $\mathcal{T}_{\text{aux}}=\left\{\mathcal{T}_{1},\dots,\mathcal{T}_{K}\right\}$ are separate MDPs that share $\mathcal{S},\mathcal{A},\mathcal{P},\rho_{0}$ and $\gamma$ with the main task $\mathcal{T}_{\text{main}}$. Each task $\mathcal{T}\in\mathcal{T}_{\text{all}}$ has an example state buffer $\mathcal{B}^{*}_{\mathcal{T}}$, where $\mathcal{T}_{\text{all}}=\mathcal{T}_{\text{aux}}\cup\left\{\mathcal{T}_{\text{main}}\right\}$ is the set of all tasks and we refer to all task-specific entities in our model as $(\cdot)_{\mathcal{T}}$.
The agent consists of multiple components for each task $\mathcal{T}$, specifically $Q_{\mathcal{T}}$, $\pi_{\mathcal{T}}$, and optionally $D_{\mathcal{T}}$. Each $\pi_{\mathcal{T}}$ maximizes the joint policy objective
$\mathcal{L}(\pi)=\sum_{\mathcal{T}\in\mathcal{T}_{\text{all}}}\mathbb{E}_{s\sim\mathcal{B}_{\text{all}},a\sim\pi_{\mathcal{T}}(\cdot|s)}\left[Q_{\mathcal{T}}(s,a)\right],$ (4)
and each $Q_{\mathcal{T}}$ minimizes the joint Bellman residual
$\mathcal{L}(Q)=\sum_{\mathcal{T}\in\mathcal{T}_{\text{all}}}\mathbb{E}_{(s,\cdot,s^{\prime})\sim\mathcal{B}}\left[(Q_{\mathcal{T}}-y_{\mathcal{T}}(s,s^{\prime}))^{2}\right]+\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}_{\mathcal{T}}}\left[(Q_{\mathcal{T}}-y_{\mathcal{T}}(s^{*},s^{*}))^{2}\right],$ (5)

### v2

# appReferences

@@ Abstract
Learning from examples of success is an appealing approach to reinforcement learning but it presents a challenging exploration problem, especially for complex or long-horizon tasks. This work introduces value-penalized auxiliary control from examples (VPACE), an algorithm that significantly improves exploration in example-based control by adding examples of simple auxiliary tasks. For instance, a manipulation task may have auxiliary examples of an object being reached for, grasped, or lifted. We show that the naïve application of scheduled auxiliary control to example-based learning can lead to value overestimation and poor performance. We resolve the problem with an above-success-level value penalty. Across both simulated and real robotic environments, we show that our approach substantially improves learning efficiency for challenging tasks, while maintaining bounded value estimates. We compare with existing approaches to example-based learning, inverse reinforcement learning, and an exploration bonus. Preliminary results also suggest that VPACE may learn more efficiently than the more common approaches of using full trajectories or true sparse rewards. Videos, code, and datasets: https://papers.starslab.ca/vpace.
I Introduction
Robotics presents a unique challenge to learning algorithms: because robotics data is costly to gather, ensuring that algorithms have high sample efficiency is crucial. In reinforcement learning and imitation learning, common approaches to improving sample efficiency incorporate a manually-defined dense reward function or demonstration trajectories of an expert completing the task, respectively. These additions, if they are available at all, can be highly biased or suboptimal. Sparse reward functions are less biased [1], but require significant exploration [2]. Sparse reward functions may also be unavailable, because they require accurate state estimation and the tuning of various parameters (e.g., how high must an object be lifted?).
@@ I Introduction
We consider another form of feedback—example states of completed tasks, sometimes referred to as example-based control (EBC) [3]. Obtaining example states can be far less laborious than designing a reward function or gathering trajectories; practitioners can gather states from a distribution that represents a completed task without consideration of how the states are reached. However, just as in the case of sparse rewards, excluding information on how goal states are reached can lead to highly inefficient learning (e.g., an example of a loaded dishwasher provides no information about the long sequence of actions required to complete the task). We aim to address the following question: Can we improve the sample efficiency of example-based control?
Figure 1: Example-based control (EBC) is inefficient due to poor exploration resulting from the inherently sparse nature of individual examples of success as a means for feedback. Auxiliary control from examples (ACE) remedies this by adding scheduled exploration of semantically meaningful auxiliary tasks, but can result in poor performance due to the unbounded value error of highly exploratory data. We combine ACE with value penalization (VP) as VPACE to efficiently learn from examples of success.
To answer this question, we propose a new example-based control method, value-penalized auxiliary control from examples (VPACE). Since hierarchical RL (HRL) has shown success in improving exploration in robotics domains [4], we leverage an HRL approach known as scheduled auxiliary control (SAC-X) [5] to improve exploration in EBC. The SAC-X framework enables practitioners to introduce a set of simple and often reusable auxiliary tasks, in addition to the main task, that help the agent to explore the environment. In this work, the full set of auxiliary tasks that we use across all main tasks are reach, grasp, lift, and release. Instead of defining main and auxiliary tasks with sparse rewards [5] or full expert trajectories [6], we define tasks with examples. For each auxiliary task there is a corresponding auxiliary policy that learns to match the set of examples. A scheduler periodically chooses and executes the different auxiliary policies, in addition to the main policy, generating a more diverse set of data to learn from.
Figure 2: Average performance across all environments and tasks studied in this work. Results are shown as an interquartile mean with five seeds per algorithm and task, and shaded regions show 95% stratified bootstrap confidence intervals [7].
We find that the naïve application of SAC-X to EBC can result in overestimated values, leading to sample inefficiency and poor performance. To remedy this, and show how SAC-X can be effectively applied to EBC, this work makes the following contributions: (i) we demonstrate that the introduction of the SAC-X framework significantly improves exploration and learning efficiency in EBC, (ii) we remedy this overestimation problem by introducing a value-penalization method that is based on the expected value of the examples, (iii) we conduct experiments across four environments with 19 simulated and two real tasks to show the improved sample efficiency and final performance of VPACE over EBC, inverse reinforcement learning, and an exploration bonus. (iv) We compare to the use of full trajectories and true sparse rewards, and observe that VPACE has higher sample efficiency than both.
II Related Work
@@
Sparse rewards are a desirable form of feedback for learning unbiased, optimal policies in reinforcement learning (RL), but they are not always obtainable, and present an immense exploration challenge on long-horizon tasks [8]. Reward shaping [9] and dense rewards can help alleviate the exploration problem in robotics [10, 11], but designing dense rewards is difficult for practitioners [12]. An alternative to manually-defined rewards is to perform inverse RL (IRL), in which a reward function is recovered from demonstrations, and a policy is learned either subsequently [13] or simultaneously via adversarial imitation learning (AIL) [14, 15, 16, 17].
@@
Like dense rewards, full trajectory demonstrations can be hard to acquire, suboptimal, or biased. Unlike IRL/AIL, in example-based control (EBC), a learning agent is only provided distributions of single successful example states. Previous EBC approaches include using generative AIL (GAIL, [16]) directly (VICE, [18]), soft actor critic (SAC, [19]) with an additional mechanism for generating extra success examples (VICE-RAQ, [20]), performing offline RL with conservative Q-learning (CQL, [21]) and a learned reward function [22], and using SAC with a classifier-based reward (RCE, [3]).
@@
All EBC methods can naturally suffer from poor exploration, given that success examples are akin to sparse rewards. Hierarchical reinforcement learning (HRL) aims to leverage multiple levels of abstraction in long-horizon tasks [23], improving exploration in RL [24, 4, 25]. Scheduled auxiliary control (SAC-X, [5]) combines a scheduler with semantically meaningful and simple auxiliary sparse rewards or auxiliary full expert trajectories (LfGP, [6, 26, 27]). Like us, [28] combined EBC with hierarchical learning, but their approach required a symbolic planner at test time, and generated very slow policies: our policies are fast and reactive, and don’t require high-level planning at test time.
Learning value functions in the off-policy setting can be challenging due to the deadly triad [29]. Regularization and clipping techniques have been applied to address various problems such as stabilizing the bootstrapping target [12, 30] and preventing overfitting and out-of-distribution samples [21, 31]. Our proposed value-penalization technique specifically targets overestimation in EBC.
TABLE I: VPACE fills a gap in literature.
Single-task Sched. Aux. Ctrl
Dense Reward RL [5]
Sparse Reward [9, 12] [5]
Dense Expert IRL [26, 6, 27]
Sparse Expert [3, 18, 22] VPACE
@@
III Example-Based Control with Value-Penalization and Auxiliary Tasks
@@
Our goal is to generate an agent that can complete a task given, as opposed to rewards or demonstrations, only final-state examples of a successfully completed task, with as few environment interactions as possible. We also assume access to final state examples of a small set of reusable auxiliary tasks. We begin by formally describing the problem setting for example-based control in Section III-A. In Section III-B, we describe how scheduled auxiliary tasks can be applied to example-based control. Finally, motivated by the increased exploration diversity of the multitask framework, we propose a new Q-estimation objective in Section III-C that leverages value penalization for improved learning stability.
III-A Problem Setting
@@
A Markov decision process (MDP) is defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,\mathcal{P},\rho_{0},\gamma\rangle$, where the sets $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action space, $\mathcal{P}$ is the state-transition environment dynamics distribution, $\rho_{0}$ is the initial state distribution, $\gamma$ is the discount factor, and the true reward $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is unknown. Actions are sampled from a stochastic policy $\pi(a|s)$. The policy $\pi$ interacts with the environment to yield experience $\left(s_{t},a_{t},s_{t+1}\right)$, generated by $s_{0}\sim\rho_{0}(\cdot),a\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$. The gathered experience $\left(s_{t},a_{t},s_{t+1}\right)$ is then stored in a buffer $\mathcal{B}$, which may be used throughout learning. For any variables $x_{t},x_{t+1}$, we may drop the subscripts and use $x,x^{\prime}$ instead when the context is clear.
@@
In this work, we focus on example-based control (EBC), a more difficult form of imitation learning where we have no access to rewards or demonstrations, but instead are given a set of example states of a completed task: $s^{*}\in\mathcal{B}^{*}$, where $\mathcal{B}^{*}\subseteq\mathcal{S}$ and $\lvert\mathcal{B}^{*}\rvert<\infty$. The goal is to (i) leverage $\mathcal{B}^{*}$ and $\mathcal{B}$ to learn or define a state-conditional reward function $\hat{R}:\mathcal{S}\to\mathbb{R}$ that satisfies $\hat{R}(s^{*})\geq\hat{R}(s)$ for all $(s^{*},s)\in\mathcal{B}^{*}\times\mathcal{B}$, and (ii) learn a policy $\hat{\pi}$ that maximizes the expected return $\hat{\pi}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\hat{R}(s_{t})\right]$.
For any policy $\pi$, we can define the value function and Q-function respectively to be $V^{\pi}(s)=\mathbb{E}_{\pi}\left[Q^{\pi}(s,a)\right]$ and $Q^{\pi}(s,a)=\hat{R}(s)+\gamma\mathbb{E}_{\mathcal{P}}\left[V^{\pi}(s^{\prime})\right]$, corresponding to the return-to-go from state $s$ (and action $a$). Temporal difference (TD) algorithms aim to estimate $V^{\pi}$ or $Q^{\pi}$ to evaluate a policy [29]. Given a reward model $\hat{R}(\cdot)$, we can say $\hat{R}(s^{*}),s^{*}\in\mathcal{B}^{*}$, indicates reward for successful states and $\hat{R}(s),s\in\mathcal{B}$, for all other states. Assuming that $s^{*}$ transitions to itself, then for policy evaluation with mean-squared error (MSE), we can write the TD targets for non-successful and successful states, $y:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$, of Q-updates as
$y(s,s^{\prime})=\hat{R}(s)+\gamma\mathbb{E}_{\pi}\left[Q(s^{\prime},a^{\prime})\right],$ (1)
$y(s^{*},s^{*})=\hat{R}(s^{*})+\gamma\mathbb{E}_{\pi}\left[Q(s^{*},a^{\prime})\right],$ (2)
where $(s,\cdot,s^{\prime})\sim\mathcal{B}$ and $s^{*}\sim\mathcal{B}^{*}$. The reward function $\hat{R}$ can be based on a learned discriminator that differentiates between $s\sim\mathcal{B}$ and $s^{*}\sim\mathcal{B}^{*}$, akin to a state-only version of adversarial imitation learning [14, 16, 17], or defined directly, such as $\hat{R}(s^{*})=1$ and $\hat{R}(s)=0$ [3, 15].
III-B Learning a Multitask Agent from Examples
We alleviate the challenging exploration problem of EBC by introducing auxiliary control from examples (ACE), an application of the scheduled auxiliary control framework [5, 6]. This hierarchical approach aims to improve exploration by learning from examples of simpler auxiliary tasks, in addition to a main task, to impose a curriculum where the completion of easier tasks provides guidance on how to complete harder ones [32]. Auxiliary tasks can be selected to have semantic meaning: in this work, the auxiliary tasks used are reach, grasp, lift, and release. The task definitions, and sometimes the examples themselves, can also be reusable between main tasks. Each task has a corresponding intention that learns from the given task-specific examples. ACE leverages a scheduler to sequentially choose and execute individual intention policies allowing for more diverse state coverage to facilitate faster learning of the main task. Crucially, because the method is off-policy, each intention can learn from data generated by any other intention, and all policy interaction data, regardless of which intention generated it, is stored in $\mathcal{B}$.
Formally, given an MDP $\mathcal{M}$, a task $\mathcal{T}$ is defined by a task-specific example buffer $\mathcal{B}^{*}_{\mathcal{T}}$. EBC methods such as RCE [3] aim to exclusively complete the main task $\mathcal{T}_{\text{main}}$, while ACE adds auxiliary tasks $\mathcal{T}_{\text{aux}}=\left\{\mathcal{T}_{1},\dots,\mathcal{T}_{K}\right\}$ during learning. We refer the set of all tasks as $\mathcal{T}_{\text{all}}=\mathcal{T}_{\text{aux}}\cup\left\{\mathcal{T}_{\text{main}}\right\}$. ACE agents are composed of two types of policies—intentions for each task and a scheduler.
III-B1 Intentions
For each task $\mathcal{T}\in\mathcal{T}_{\text{all}}$, the corresponding intention consists of a task-specific policy $\pi_{\mathcal{T}}$, Q-function $Q_{\mathcal{T}}$, and state-conditioned reward $\hat{R}_{\mathcal{T}}$. ACE optimizes the task-specific policies by maximizing the policy optimization objective
$\mathcal{L}(\pi;\mathcal{T})=\mathbb{E}_{\mathcal{B},\pi_{\mathcal{T}}}\left[Q_{\mathcal{T}}(s,a)\right].$ (3)
The task-specific Q-functions are optimized via minimization of the Bellman residual
$\begin{split}\mathcal{L}(Q;\mathcal{T})=&\mathbb{E}_{\mathcal{B},\pi_{\mathcal{T}}}\left[(Q_{\mathcal{T}}(s,a)-y_{\mathcal{T}}(s,s^{\prime}))^{2}\right]\\ &+\mathbb{E}_{\mathcal{B}^{*}_{\mathcal{T}},\pi_{\mathcal{T}}}\left[(Q_{\mathcal{T}}(s^{*},a)-y_{\mathcal{T}}(s^{*},s^{*}))^{2}\right],\end{split}$ (4)